5. A student takes a multiple-choice exam where each question has 5 possible answers. He works a question correctly if he knows the answer, otherwise he guesses at random. Suppose he knows the answer to 80% of the questions. (a) What is the probability that on a question chosen at random the student gets the correct (b) Given that the student gets the correct answer to this question, what is the probability answer? that he actually knew the answer?
Answer #1

a)P(gets correct answer)=P(knows and gets correct answer)+P(guess and gets correct answer)

=0.8*1+0.2*1/5 =0.84

b)P(knew the answer given gets correct answer)=P(knows and gets correct answer)/P(gets correct answer)

=0.8*1/0.84 = 0.9524